OpenStudy (anonymous):

I cannot figure out this logarithm! 8^(2x+1)=16^(1-x) I rewrote the coefficients to have the same base...2^3 and 2^4, what do i do next?? 2^3(2x+1)=2^4(1-x) do i factor both sides or whatt?? :(

OpenStudy (anonymous):

\[\Large 8^{2x + 1} = 16^{1 - x}\] \[\Large 2^{6x + 3} = 2^{4 - 4x}\] Divide both sides by 2^(4 -4x) \[\Large \frac{2^{6x+3}}{2^{4 - 4x}} = 1\] \[\Large 2^{6x + 3 - (4 - 4x)} = 2^{10x - 1} = 1\] \[\Large 2^{10x} = 2^1\] \[10x = 1\] \[x = \frac{1}{10}\]

OpenStudy (dumbcow):

since they have same base .... take log (base 2) of both sides leaving --> 3(2x+1) = 4(1-x)

OpenStudy (anonymous):

^ better way than mine

OpenStudy (anonymous):

i just had a serious math meltdown. thank you both very much for clearing that up!!!